OpenStudy (anonymous):

If matrix A has an inverse A^-1, use equation 2: Av = lambda*v, to show that A^-1 has the same eigenvectors as A. Determine a relationship between the eigenvalues of A and A^-1. Illustrate with a suitable example.

OpenStudy (phi):

what happens if you multiply both sides by A^-1 ?

OpenStudy (anonymous):

Then eventually you'd get A^-1 = 1/lambda, but there was another problem like this and this was the answer to it, but this one is a little different, so u know what to do here?

OpenStudy (phi):

you start with \[ Av = \lambda v\] multiply both sides by A inverse: \[ A^{-1}Av = A^{-1} \lambda v\] what do you get on the left side? what do you get on the right side?

OpenStudy (anonymous):

On the left side, you get just v and on the right side you get A^-1 lambda v

OpenStudy (phi):

yes, and remember A^-1 is a matrix, lambda is a scalar (a simple number) and v is a vector. when you multiply a number times a matrix, you can switch the order, so you can write it as: \[ v= \lambda A^{-1}v \] now divide both sides by lambda (assuming it is not zero) \[ \frac{1}{\lambda} v = A^{-1}v \] or \[ A^{-1}v = \frac{1}{\lambda} v \] Does that look familiar? Matrix times a vector = a number times that same vector

OpenStudy (anonymous):

So then does that answer our question? And what would be a suitable example like it tells us to do?

OpenStudy (phi):

if you see Av = lambda*v then you know that v is an eigenvector of matrix A and lambda is an eigenvalue associated with eigenvector v So what does \[ A^{-1}v = \frac{1}{\lambda} v \] tell you?

OpenStudy (phi):

v is the ____ of matrix A^-1 and 1/lambda is the _____ associated with v

OpenStudy (anonymous):

So v is the eigenvector and 1/lambda is the eigenvalue? Is this right?

OpenStudy (phi):

yes. This shows that A and its inverse have the same eigenvectors, and their eigenvalues are the reciprocal (1/lambda and lambda)

OpenStudy (anonymous):

So the answer would be that A and A^-1 are reciprocals of each other, would this be our answer?

OpenStudy (phi):

So the answer would be that **the eigenvalues of *** A and A^-1 are reciprocals of each other, would this be our answer?

OpenStudy (phi):

is that what you mean?

OpenStudy (anonymous):

yea